A sample of 6.44g of naphthalene (C10H8) is dissolved in 80.1g of benzene (C6H6). The density of naphthalene is 1.14g/mL and the density of benzene is 0.88g/mL.
1
Expert's answer
2020-02-28T12:12:30-0500

"\rho"=(m1+m2)/(V1+V2)=(6.44+80.1)/(80.1/0.88+6.44/1.14)=0.90 g/mL